### v1

@@ IV Conclusion
$\vspace{-5cm}\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&\frac{1}{3}&-\frac{1}{3}&\frac{1}{3\sqrt{2}}&-\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&-\frac{1}{2}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&\frac{2}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{6}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{5}{2})\end{pmatrix}$ (48)
@@
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\mathcal{C}_{T(P)}\left(\begin{array}[]{cccccc}-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ 0&0&0&0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}\\ 0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ 0&0&\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&0&0&0\\ 0&0&0&0&0&\frac{1}{\sqrt{2}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ 0&0&0&0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}\\ 0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ 0&0&\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&0&0&0\\ 0&0&0&0&0&\frac{1}{\sqrt{2}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ \end{pmatrix}$ (77)

### v2

@@ IV Conclusion
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}-\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&\frac{\sqrt{2}}{3}&-\frac{1}{3}&\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&-\frac{1}{\sqrt{5}}\\ 0&0&0&-\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&-\frac{2}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{5}{2})\end{pmatrix}$ (54)
@@
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\mathcal{C}_{T(P)}\left(\begin{array}[]{cccccc}\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ 0&0&0&0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}\\ 0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ 0&0&-\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&\frac{\sqrt{2}}{3}&\frac{1}{3}\\ 0&\frac{\sqrt{2}}{3}&\frac{1}{3}&0&0&0\\ 0&0&0&0&0&-\frac{1}{\sqrt{2}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ 0&0&0&0&-\frac{2}{3}&\frac{1}{3\sqrt{2}}\\ 0&-\frac{2}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ -\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ \frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ 0&0&\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&0&0&0\\ 0&0&0&0&0&\frac{1}{\sqrt{2}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ \end{pmatrix}$ (83)